A particular fruit's weights are normally distributed, with a mean of 492 grams and a standard deviation of 24 grams.

If you pick 4 fruits at random, then 8% of the time, their mean weight will be greater than how many grams?

Enter your answer as an equation or as a number to the nearest gram.
Answer #1
mean μ= 492
standard deviation σ= 24.0000
sample size       =n= 4
std error=σ=σ/√n= 12

as top 8% values are at 92th percecentile for which critical z =1.41

therefore corresponding value=mean+z*std deviation= ~ 509 gram